OpenStudy (anonymous):

Given the following functions f(x) and g(x), solve f[g(7)] and select the correct answer below: f(x) = 4x + 21 g(x) = 2x + 2 16 49 85 784

OpenStudy (anonymous):

@Icedragon50 anything

OpenStudy (anonymous):

??

Vocaloid (vocaloid):

g(7) = ?

OpenStudy (anonymous):

7g

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

@Vocaloid

OpenStudy (haseeb96):

f(g(x)) = f(2x+2) f(g(7) = f(2(7) + 2) = f(14+2) = f(16) =4(16) +21 now solve it yourself and get your answer

OpenStudy (haseeb96):

@drkid15 did you get it?

OpenStudy (anonymous):

85

OpenStudy (anonymous):

@Haseeb96

OpenStudy (anonymous):

hello @Haseeb96

OpenStudy (haseeb96):

correct answer @drkid15 good job